OpenStudy (anonymous):

x=the square root of 4x+12

OpenStudy (atlas):

x=-6,4

OpenStudy (anonymous):

\[Solve the equation: x=\sqrt{4x+12}\]

OpenStudy (atlas):

x2 -4x-12=0 x2-6x+2x-12=0 x(x-6)+2(x-6)=0 (x+2)(x-6)=0 x=-2,6

OpenStudy (anonymous):

x=6,-2